OpenStudy (anonymous):

how do i solve u=3x+1/2x+5 for x... show work pls

myininaya (myininaya):

\[u=\frac{3x+1}{2x+5}\] Multiply (2x+5) on both sides.

OpenStudy (anonymous):

got that far

myininaya (myininaya):

\[(2x+5)u=3x+1\] Distribute! \[2x u +5 u =3x +1\]

myininaya (myininaya):

put terms with u on one side and terms without u on the other side

OpenStudy (anonymous):

alright, then haha?

myininaya (myininaya):

\[2 x u-3 x= 1- 5u\] sorry i wrote that backwards terms with x on one side and terms without x on the other

myininaya (myininaya):

now factor \[x(2 u -3) =1- 5u\]

myininaya (myininaya):

\[x=\frac{1 - 5u}{2u -3}\]

OpenStudy (anonymous):

got it. thank you :D